OpenStudy (anonymous):

Help! 133 1/2 +84 1/2+ 124 5/8 + 128 1/4=

hero (hero):

Go find a calculator

OpenStudy (anonymous):

(133 1/2) + (84 1/2) + (124 5/8) + (128 1/4) = 470.87500

hero (hero):

yeah, the answer should be in fraction form

hero (hero):

not decimal

OpenStudy (anonymous):

ok, I have reduced the equation to 133/2+84/2+31*5/2+64/2 = 436/2

hero (hero):

I think you changed it by doing that

hero (hero):

133 1/2 doesn not equal 133/2

OpenStudy (anonymous):

why?

hero (hero):

133 1/2 = 267/2

hero (hero):

Think mixed numbers here

hero (hero):

Anyways, it's 470 7/8

OpenStudy (anonymous):

ok i didn't know they were mixed fraction . I was treating them as fractions.

OpenStudy (anonymous):

yes, @hero is right , 267/2+169/2+997/8+513/4 => 218 + 997/8+513/4 taking lcm we get 1744+997+1026/8 = 3767/8 = 470 7/8.